OpenStudy (anonymous):

(1/4-1/5i)(4/5+4/5i) MULTIPLY!

OpenStudy (anonymous):

\[(a+bi)(c+di)=(ac-bd)+(ad+bc)i\]

OpenStudy (anonymous):

How is it with the i (imaginary number)

OpenStudy (anonymous):

1/4 * 4/5 = 17/20 ?

OpenStudy (anonymous):

17/20 + 17/20i - 16/20i - 16/20 ?